A20. Emmanuel is given a multiple-choice exam with ten questions and each question with five possible answers. He decided to guess randomly for each question. What is the probability that he will get at least six questions correct?
1
Expert's answer
2020-06-04T21:20:58-0400

Probability of question correct "= \frac{1}{5}"


Probability of question Incorrect "=\frac {4}{5}"


Probability get at least six questions correct


"=_{10}C_{6} (\frac{1}{5})^6(\frac {4}{5})^4 +_{10}C_{7} (\frac{1}{5})^7(\frac {4}{5})^3+_{10}C_{8} (\frac{1}{5})^8(\frac {4}{5})^2+_{10}C_{9} (\frac{1}{5})^9(\frac {4}{5})^1+_{10}C_{10} (\frac{1}{5})^{10} (\frac {4}{5})^0"


"= 6.3739 \times 10 ^{-3}"